### v4

# OpenFly: A Versatile Toolchain and Large-scale Benchmark for Aerial Vision-Language Navigation

Abstract
Vision-Language Navigation (VLN) aims to guide agents through an environment by leveraging both language instructions and visual cues, playing a pivotal role in embodied AI. Indoor VLN has been extensively studied, whereas outdoor aerial VLN remains underexplored. The potential reason is that outdoor aerial view encompasses vast areas, making data collection more challenging, which results in a lack of benchmarks. To address this problem, we propose OpenFly, a platform comprising a versatile toolchain and large-scale benchmark for aerial VLN. Firstly, we develop a highly automated toolchain for data collection, enabling automatic point cloud acquisition, scene semantic segmentation, flight trajectory creation, and instruction generation. Secondly, based on the toolchain, we construct a large-scale aerial VLN dataset with 100k trajectories, covering diverse heights and lengths across 18 scenes. The corresponding visual data are generated using various rendering engines and advanced techniques, including Unreal Engine, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS). All data exhibit high visual quality. Particularly, 3D GS supports real-to-sim rendering, further enhancing the realism of the dataset. Thirdly, we propose OpenFly-Agent, a keyframe-aware VLN model, which takes language instructions, current observations, and historical keyframes as input, and outputs flight actions directly. Extensive analyses and experiments are conducted, showcasing the superiority of our OpenFly platform and OpenFly-Agent. The toolchain, dataset, and codes will be open-sourced.
Figure 1: Overview of OpenFly. This work consists of an automatic toolchain for data generation, a large-scale aerial VLN dataset comprising 100K trajectories and instructions, and a keyframe-aware VLN model emphasizing key observations.
I Introduction
Figure 2: Framework of the automatic data generation platform. Various rendering engines and simulators are first integrated, providing diverse high-quality scenes. Built on these, several interfaces and tools are developed, enabling automated generation of trajectories and instructions.
Embodied AI has drawn increasing research interest, where vision-language navigation (VLN) is a key task, aiming to navigate agents to a target location according to linguistic instructions and visual observations. Recently, many VLN datasets have been established, e.g., TouchDown  [7], REVERIE [37], R2R [1], RxR [25], CVDN [45], VLN-CE [23], and LANI [33], which significantly facilitate a series of VLN methods  [30, 9, 40, 15, 17, 52, 51, 10, 56]. However, all these works focus on indoor or ground-based agents, overlooking unmanned aerial vehicles (UAVs) that play an important role in aerial photography, rescue operations, and cargo transportation.
Most recently, AerialVLN [29] and OpenUAV [47] have made significant strides by leveraging UAV simulators to bridge gaps in aerial VLN datasets, thus greatly advancing the development of this field. However, several challenges remain that need to be addressed:
  • Insufficient data diversity. Existing methods rely on AirSim and Unreal Engine (UE) to control UAVs, limiting them to digital assets compatible with these platforms, reducing the diversity of available data and restricting the potential use of more photorealistic sources.
  • Inefficient data collection. The process of generating trajectories relies on pilots operating UAVs in simulators, followed by manual annotation to create language instructions. The entire process is labor-intensive, time-consuming, and difficult to scale.
  • Inadequate data scale. Current datasets for aerial VLN remain relatively small, with only around 10k trajectories, which is far behind embodied manipulation datasets. For example, Open X-Embodiment [36] has collected 1M episodes of manipulation data, which has significantly accelerated the development of vision-language-action (VLA) models [38, 54, 35, 55].
To address these issues, we propose OpenFly, a comprehensive platform consisting of a versatile toolchain and a large-scale benchmark for the aerial VLN task. To enhance data diversity, the platform is established on various widely-used rendering engines and advanced techniques, i.e., UE, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS). This enables us to utilize a wide range of assets as shown in Fig. 1. In particular, we use UAVs for automated patrols to capture numerous real-world images and integrate 3D GS technology into our platform, through which we reconstruct realistic 3D scenes. This empowers our platform to support a real-to-sim method, enhancing the realism of data. To improve the efficiency of data collection, we develop a versatile toolchain for automated aerial VLN data generation, as depicted in Fig. 2. Specifically, point cloud acquisition is first conducted to capture the 3D occupancy of a scene. Next, scene semantic segmentation is performed to identify and select landmarks as waypoints along the flight trajectories. Building on these tools, trajectory generation is then carried out, taking landmarks and point clouds as input, using predefined flight actions as basic units, and automatically searching for a collision-free trajectory. Finally, we feed the trajectories and corresponding UAV-egocentric images into a vision-language-model (VLM), i.e., GPT-4o, to generate linguistic instructions. The entire pipeline is highly automated, reducing the reliance on UAV pilots and annotators. To collect a large-scale dataset, we meticulously collected 18 high-quality scenes, generating various trajectories of differing heights and lengths. Benefitting from our toolchain, we are able to quickly construct a dataset of 100k samples, significantly larger than existing datasets.
Besides, we propose OpenFly-Agent, a keyframe-aware aerial VLN model trained on the OpenFly dataset. It is developed based on a VLA model OpenVLA [22] and incorporates an adaptive frame-level token-sampling mechanism to mitigate visual distraction caused by redundant historical observations. Note that this sampling mechanism is crucial for the visual encoding of aerial VLN, as UAVs fly quickly and observations change rapidly. Extensive experiments are conducted on the OpenFly dataset to evaluate numerous methods, establishing a comprehensive benchmark for the aerial VLN tasks.
Overall, our contributions can be summarized as follows:
  • We build OpenFly on multiple rendering engines and develop a versatile toolchain, enabling the automatic generation of data with both great diversity and efficiency.
  • We have constructed a novel aerial VLN benchmark comprising 100k trajectories across 18 high-quality scenes. To the best of our knowledge, this is the largest aerial VLN benchmark to date.
  • We propose OpenFly-Agent, a keyframe-aware VLN model. Extensive experiments demonstrate its superior performance compared to other methods.
II Related Works
II-A Simulators for Embodied AI
In recent years, many simulators related to embodied AI have been developed, e.g., MuJoCo [46], Habitat [42], Pybullet [11], Matterport3D [6], OpenAI Gym [4], and Issac Gym [32]. They allow researchers to learn and validate algorithms in safe virtual environments, significantly advancing the development of embodied intelligence. However, most of them are designed for indoor robotic manipulation and navigation, making them unsuitable as simulators for aerial VLN tasks. Gazebo and AirSim [44] are two popular UAV simulators, while Gazebo suffers from low rendering quality. Combining AirSim and UE is a potential solution to construct aerial VLN scenes. Unfortunately, the official maintenance and updates for AirSim have been discontinued, resulting in a limited availability of compatible assets.
II-B Vision-Language Navigation Datasets
Numerous datasets have been proposed to accelerate the VLN task. R2R [1] focuses on evaluating agents in unseen buildings and provides discrete navigation options. RxR [25] is a densely annotated VLN dataset, supporting multilingual instructions. TouchDown [7] and REVERIE [37] have each contributed a dataset from real-life environments, which requires a ground-based agent to follow instructions and find a target object. CVDN [45] presents a cooperative VLN dataset where agents can access the dialog history of human cooperation for inference. All the above datasets are graph-based where navigable points are predefined. LANI [33] and VLN-CE [23] propose the VLN task in continuous outdoor/indoor environments, enabling agents to move freely to any unobstructed point. Recently, a few works have tried to construct VLN datasets for aerial space. ANDH [12] establishes a dialogue-based aerial VLN dataset with bird-view images. CityNav [26] builds on the point cloud data from SensatUrban [18] and linguistic annotations from CityRefer [34], which requires a real-world 2D map to help locate specific landmarks in the instruction. AerialVLN [29] and OpenUAV [47] integrate AirSim and UE to create VLN scenes where pilots can control UAVs to generate various trajectories.
II-C Vision-Language Navigation Methods
VLN methods require agents to understand and complete language instructions according to visual observations. In graph-based works [31, 49, 20, 13], agents move among predefined graph nodes. Recently, LLM-driven works [58, 57, 8] use the reasoning ability of large language models to infer the next node. Although great progress has been achieved, these methods are not applicable in real unknown environments. In contrast,  [41, 19, 24, 56] explore VLN in continuous environments for more realistic applications. More recently, a few works attempt to address the problem of aerial VLN based on UAVs, among which AerialVLN [29] proposes a lookahead guidance method to generate more reasonable ground-truth trajectories during training, and STMR [14] designs a matrix representation to improve the spatial reasoning of LLMs.
III OpenFly Data Generation Platform
Figure 3: High-quality examples from different rendering engines and techniques, including several large cities such as Shanghai, Guangzhou, Los Angeles, Osaka, and etc., cover an area of over a hundred square kilometers in total. 3D GS provides five large campus scenes, further enhancing the diversity and realism of the data.
In this section, we first describe several basic simulators and data resources, and then present the developed toolchain. The framework of the whole automatic data generation platform is illustrated in Fig. 2.
III-A Basic Simulators and Data Resources
To collect a wide range of high-quality and realistic simulation data, we source the dataset from multiple rendering engines integrated with various simulators. Fig. 3 showcases several examples obtained from these rendering engines/techniques.
Unreal Engine + AirSim/UnrealCV. UE is a rendering engine capable of providing highly realistic interactive virtual environments. This platform has undergone five iterations, and each version features comprehensive and high-quality digital assets. In UE5, we meticulously select an official sample project named ‘City Sample’, which provides us with a large urban scene covering $25.3km^{2}$ and a smaller one covering $2.7km^{2}$. These scenes include a variety of assets such as buildings, streets, traffic lights, vehicles, and pedestrians. Besides, in UE4, we prepare six more high-quality scenes. Specifically, there are two large scenes showcasing the central urban areas of Shanghai and Guangzhou, covering areas of $30.88km^{2}$ and $58.56km^{2}$, respectively. The remaining four scenes are selected from AerialVLN [29]. They have smaller areas for totally about $26.64km^{2}$. These scenes encompass a wide range of architectural styles, including both Chinese and Western influences, as well as classical and modern designs. Additionally, the UE4 engine allows us to make adjustments in scene time to achieve different appearances of scenes under varying lighting conditions.
Airsim is an open-source simulator, which provides highly realistic simulated environments for UAVs and cars. We integrate the AirSim plugin into UE4 to obtain image data easily from the perspective of a UAV. Since AirSim does not support UE5 and stopped updating in 2022, we use the UnrealCV [53] plugin as an alternative for image acquisition in UE5. To realize a highly efficient data collection in simulated scenes, we modify the UE5 project to a C++ project, integrate the UnrealCV plugin, and package executables for multiple systems like Windows and Linux.
GTA V + Script Hook V. GTA V is an open-world game that is frequently used by computer vision researchers due to its highly realistic and dynamic virtual environment. The game features a meticulously crafted cityscape modeled after Los Angeles, encompassing various buildings and locations such as skyscrapers, gas stations, parks, and plazas, along with dynamic traffic flows and changes in lighting and shadows.
Script Hook V is a third-party library with the interface to GTA V’s native script functions. With the help of Script Hook V, we build an efficient and stable interface, which receives the pose information and returns accurate RGB images and lidar data. From the interface, we can control a virtual agent to collect the required data in an arbitrary pose and angle in the game.
Google Earth + Google Earth Studio. Google Earth is a virtual globe software, which builds a 3D earth model by integrating satellite imagery, aerial photographs, and Geographic Information System (GIS) data. From this engine, we select four urban scenes covering a total area of $53.60km^{2}$, i.e., Berkeley, primarily consisting of traditional neighborhoods; Osaka, which features a mix of skyscrapers and historic buildings; and two areas with numerous landmarks: Washington, D.C., and St. Louis.
Google Earth Studio is a web-based animation and video production tool that allows us to create keyframes and set camera target points on the 2D and 3D maps of Google Earth. Using this functionality, we can quickly generate customized tour videos by selecting specific routes and angles. In order to efficiently plan the route, we develop a function that automatically draws the flight trajectory in Google Earth Studio according to the selected area and predefined photo interval.
3D Gaussian Splatting + SIBR viewers. As a highly realistic reconstruction method, hierarchical 3D GS  [21] employs a hierarchical training and display architecture, making it particularly suitable for rendering large-scale areas. Due to these features, we use this method to reconstruct and render multiple real scenes. We utilize the DJI M30T drone as the data collection device, which offers an automated oblique photography mode, enabling us to capture a large area of real-world data with minimal manpower. Practically, we gathered data from five campuses across three universities, which are East China University of Science and Technology, Northwestern Polytechnical University, and Shanghai Jiao Tong University (referred to as ECUST, NWPU, and SJTU). These campus scenes include various types and styles of landmarks, such as libraries, bell towers, waterways, lakes, playgrounds, construction sites, and lawns. The detailed information for the five campuses is presented in Table I. More details of the 3D GS data collection can be found in our supplementary material.
SIBR [3] viewers is a rendering tool designed for the 3D GS project, enabling visualization of a scene from arbitrary viewpoints. The tool supports high-frame-rate scene rendering and provides various interactive modes for navigation. Building upon SIBR viewers, we developed an HTTP RESTful API that generates RGB images from arbitrary poses, simulating a UAV’s perspective.
TABLE I: Different 3D GS Scenes
Campus Name Images Area
ECUST (Fengxian Campus) 12008 $1.06km^{2}$
NWPU (Youyi Campus) 4648 $0.8km^{2}$
NWPU (Changan Campus) 23798 $2.6km^{2}$
SJTU (Minghang-East Zone) 20934 $1.72km^{2}$
SJTU (Minghang-West Zone) 9536 $0.95km^{2}$
III-B Toolchain for Automatic Data Colleciton
To achieve automatic data generation, we integrate the above five simulators and design three unified interfaces, i.e., the agent movement interface, the lidar data acquisition interface, and the image acquisition interface, allowing an agent to interact with any scene. Based on these interfaces, we further develop a toolchain, including 3D point cloud acquisition, scene semantic segmentation, automatic trajectory generation, and instruction generation. The framework of the whole data generation platform is illustrated in Fig. 2, with details of these interfaces and tools elaborated below.
Unified Interfaces. 1) Agent Movement Interface: We design a CoorTrans module, which implements a customized pose transformation matrix and scaling function to unify all simulator coordinate systems into a meter-based FLU (Front-Left-Up) convention. This interface enables precise agent positioning among regular scenes, point clouds, and scene segmentations, ensuring consistency and facilitating automatic trajectory generation. 2) Lidar Data Acquisition: For different simulators, point cloud data is acquired through different methods, including lidar sensor collection, depth map back-projection, and image feature matching. We develop a unified interface to integrate these methods and leverage the proposed CoorTrans module to align all data to the same FLU coordinate system. 3) Image Acquisition Interface: We integrate HTTP RESTful and TCP/IP protocols to form a unified image request interface, allowing image data to be obtained from any location with flexible resolutions and agent viewpoints.
3D Point Cloud Acquisition. For different simulators, we provide two methods to reconstruct the point cloud map of an entire scene. 1) Rasterized Sampling Reconstruction: For the UE5 + UnrealCV simulator, the MatrixCity [28] project offers a convenient rasterized sampling solution. We use the aforementioned lidar data acquisition interface to obtain the local point cloud at the sampling points. Since these data are already aligned within the same coordinate system, the point cloud map of the entire scene can be constructed by simply stitching local point clouds. For the UE4 + AirSim and GTA V simulators, we customize rasterized sampling points at appropriate resolutions, and perform sampling and reconstruction using the agent movement and lidar data acquisition interfaces. 2) Image-based Sparse Reconstruction: In 3D GS, the scene reconstruction process begins with the open-source COLMAP [43] framework, which geneoverlearates a sparse point cloud from input images. We directly export and use the point clouds obtained from this step.
Scene Semantic Segmentation. Vision-and-Language Navigation (VLN) requires meaningful landmarks as navigation targets. We perform semantic segmentation on four types of simulation scenes using the following three methods. 1) 3D Scene Understanding: A sequence of top-down views of the scene is captured in a rasterized format. We then use Octree graph [50] to extract 2D mask proposals, which are subsequently projected into the 3D point cloud space to generate semantic 3D segments. 2) Point Cloud Projection and Contour Extraction: We first acquire the point cloud of a scene, then project the voxelized point cloud onto a projection plane slightly above the ground. Using OpenCV, we perform a series of operations on the projected image, including binarization, erosion, and contour extraction, to obtain multiple instances along with their 2D coordinates. For each instance, the maximum height of the points within its neighborhood is used as the final height. This method provides a more computationally efficient yet coarser segmentation compared to the first approach, allowing users to choose based on their requirements. 3) Manual Annotation: When the point cloud quality of a scene is low or finer segmentation is required, we provide a method for annotating instances in the point cloud space by mouse clicks, based on ROS2 and RVIZ2. Users can annotate instances directly in the point cloud space using mouse clicks to define landmarks of interest for the task. This method is applicable to four simulators, i.e., UE + UnrealCV, UE + Airsim, GTA V, and 3D GS.
Automatic Trajectory Generation. Leveraging the aforementioned point cloud map and segmentation tools, OpenFly can generate trajectories using the following two methods. 1) Path search based on customized action space: First, a global hash voxel map $M_{global}$ and a bird’s eye view (BEV) occupancy map $M_{bev}$ are constructed from the scene point cloud $P$. Second, the flight altitude is randomly selected within the user-defined height range, and landmarks that are not lower than the height threshold $H_{\tau}$ are chosen as targets. A starting point is selected within the distance range $[r,R]$ from the landmark, ensuring that it is not occupied in both $M_{global}$ and $M_{bev}$. Then, a point on the line connecting the starting point and the landmark, which is close to the landmark and unoccupied in $M_{bev}$, is chosen as the endpoint. Third, A collision-free trajectory from the starting point to the endpoint is generated using the A*  [16] pathfinding algorithm, where the granularity of exploration step size and direction can be adjusted according to the action space. Besides, by repeatedly selecting the endpoint as the new starting point, complex trajectories can be generated. Finally, utilizing OpenFly’s interface, images corresponding to the trajectory points can be obtained. 2) Path search based on grid: Google Map data does not allow image retrieval at arbitrary poses in the space. Thus we rasterize a pre-selected area and collect images from each grid point in all possible orientations. Starting and ending points are chosen within the grid points to generate trajectories. Corresponding images for these trajectory points are then selected from the pre-collected image set.
Automatic Instruction Generation. Previous research has predominantly relied on manual annotation to generate trajectory instruction, which is not only costly but also limits the scalability of datasets. To address this issue, we propose a highly automated language instruction generation method based on VLMs, e.g., GPT-4o. A straightforward method would be to submit all images to VLMs to analyze the trajectory and generate instructions. However, using all images introduces significant computational overhead and leads to information redundancy. For example, the ‘Forward’ action usually occupies a larger proportion of a flight trajectory, with ‘Turn Left/Turn Right’ or ‘Ascend/Descend’ actions taken when encountering key landmarks.
Based on the above findings, we split the complete trajectory into multiple sub-trajectories based on the occurrence of non-consistent actions, extracting key actions and images for processing and subsequent integration. Notably, slight angle adjustments often occur during flight to change the direction subtly, and a ‘slightly Turn Left/Right’ will be merged with subsequent ‘Forward’ actions. Specifically, suppose that the trajectory action sequence is [1, 1, 2, 1, 1, 2, 2, 1, 0], where 1, 2, and 0 denote ‘Forward’, ‘Turn Left’, and ‘Stop’, respectively. This trajectory would be split into four sub-trajectories, i.e., [1, 1], [2, 1, 1], [2, 2], and [1, 0]. The second sub-trajectory involves ‘slightly turn left’ and ‘move forward’. We submit the action sequence and the last captured image of each sub-trajectory to a VLM to generate descriptions of both action and landmarks. The sub-instructions are obtained similar to the following format:
  • ‘Move forward’ to ‘Landmark 1’.
  • ‘Slightly turn left and move forward’ to ‘Landmark 2’.
  • ‘Turn left’ towards ‘Landmark 3’.
  • ‘Move forward’ to ‘Landmark 4’.
These sub-instructions are then processed by a VLM/LLM again, where they are integrated into coherent and complete instructions. The detailed prompt used for the VLM, along with the complete responses, is provided in the supplementary material.
III-C Quality Control.
Data Filter. During data collection, it is inevitable that some damaged or low-quality data will be generated. We clean the data in the following situations. 1) We remove damaged images that are produced in generation or transmission. 2) We find that UAVs sometimes appear to pass through the tree models. Therefore, we exclude the trajectories where the altitude is lower than that of the trees. 3) We believe that extremely short or long trajectories are not conducive to model training. Thus, we remove these trajectories, specifically those with fewer than 2 or more than 150 actions.
Instruction Refinement. A known challenge of instruction generation is VLMs’ hallucinations. During the previous instruction generation process, sometimes the same landmark appears across several frames. This results in a VLM generating similar captions for the repeated observations of a landmark, increasing the complexity of the final instruction and introducing ambiguity due to duplication.
To mitigate this challenge, we utilize the NLTK library  [2] to simplify the instruction by detecting and merging similar descriptions. Specifically, a syntactic parse tree is first constructed to extract all landmark captions using a rule-based approach. Then, a sentence-transformer model is employed to encode the extracted landmark captions into embedding vectors. Their similarities are computed with dot product, and high-similarity captions are then identified and replaced with referential pronouns (e.g., “it,” “there,” etc.). For example, a generated instruction with redundant information is $\cdots$ make a left turn toward a medium-sized beige building marked by a signboard reading CHARLIE’S CHOCOLATE. Continue heading straight, passing a medium-sized gray building with a prominent rooftop billboard displaying Charlie’s Chocolate $\cdots$”. After simplification, a more concise sentence is obtained, i.e.$\cdots$ make a left turn toward a medium-sized beige building marked by a signboard reading CHARLIE’S CHOCOLATE. Continue heading straight, passing it $\cdots$”, demonstrating the effectiveness of this post-processing technique.
At the same time, we built a data inspection platform to provide instructions, action sequences, and corresponding images to the examiners. If the instructions and trajectories align, they are considered qualified. We randomly select 3K samples from the entire dataset according to data distribution in Sec. V-A. After manually inspecting these samples, we find that the qualification rate reaches 91%.
IV Dataset Analysis
TABLE II: Comparison of different VLN datasets. Above the middle dividing line lies the ground-based datasets, while below is the aerial VLN datasets. $N_{traj}$: the number of total trajectories. $N_{vocab}$: vocabulary size. Path Len: the average length of trajectories, measured in meters. Intr Len: the average length of instructions. $N_{act}$: the average number of actions per trajectory.
Dataset $N_{traj}$ $N_{vocab}$ Path Len. Intr Len. Action Space $N_{act}$ Environment R2R [1] 7189 3.1K 10.0 29 graph-based 5 Matterport3D RxR [25] 13992 7.0K 14.9 129 graph-based 8 Matterport3D REVERIE [37] 7000 1.6K 10.0 18 graph-based 5 Matterport3D CVDN [45] 7415 4.4K 25.0 34 graph-based 7 Matterport3D TouchDown [7] 9326 5.0K 313.9 90 graph-based 35 Google Street View VLN-CE [23] 4475 4.3K 11.1 19 2 DoF 56 Matterport3D LANI [33] 6000 2.3K 17.3 57 2 DoF 116 CHALET ANDH [12] 6269 3.3K 144.7 89 3 DoF 7 xView AerialVLN [29] 8446 4.5K 661.8 83 4 DoF 204 AirSim + UE CityNav [26] 32637 6.6K 545 26 4 DoF - SensatUrban OpenUAV [47] 12149 10.8K 255 104 6 DoF 264 AirSim + UE Ours 100K 15.6K 99.1 59 4 DoF 35 AirSim + UE, GTA5 + Script Hook V, Google Earth Studio, 3D GS + SIBR viewers
(a) Difficulty level distribution.
(b) Length-height distribution.
(c) Action distribution with 1 type of ‘Forward’.
(d) Action distribution with 3 types of ‘Forward’.
Figure 4: Statistical analysis of trajectories.
IV-A Overview
Using our toolchain, we collect 100k trajectories from 18 scenes, along with corresponding image sequences and language instructions. During the data generation process, we set a minimum motion step size of 3 meters to produce more granular trajectories. The details of our and previous VLN datasets are listed in Table. II, from which we can see that our dataset features a significantly larger number of trajectories and a more extensive vocabulary, as well as greater environmental diversity. In contrast, our average trajectory length and instruction length are relatively short. This is intentional, as we believe” short- and medium-range instructions are actually more in line with the usage habits of human users. This might be more beneficial for the aerial VLN field.
IV-B Trajectory Analysis
In addition to a rich variety of scenes, we also strive for diversity in the difficulty level, length, and height of the trajectory data. Based on the number of actions in one trajectory, we classify trajectories with fewer than 30 actions as ‘Easy’, those with the number of actions ranging from 30 to 60 as ‘Moderate’, and those with more than 60 actions as ‘Hard’. Fig. 4(a) shows the corresponding difficulty level distribution. Besides, compared with ground-based VLN, the aerial VLN task has more motion dimensions. Therefore, we set different trajectory lengths and flight heights to obtain rich data. Fig. 4(b) exhibits the distribution of these data, with their lengths ranging from 0 to 300 meters, and the heights ranging from 0 to 150 meters.
In the aerial VLN tasks of large-scale outdoor scenes, the proportion of moving forward is naturally higher than that of making adjustments in direction and altitude, as shown in Fig. 4(c). However, this highly unbalanced action distribution might cause the VLN model to overfit to the dominant action. To alleviate this problem, we divide the ‘Forward’ action into three granularities, i.e., 3m, 6m, and 9m. In the ground-truth trajectories, three consecutive ‘Forward’ actions will be combined into one ‘9m Forward’ action. At the end of a straight-moving trajectory, if the remaining distance is less than 9m, it will be combined into a ‘6m Forward’ action, or remain as a ‘3m Forward’ action. Fig. 4(d) presents the action distribution after this action merging process.
(a) Train set distribution.
(b) Test set distribution.
Figure 5: The distribution of the data volume in different scenes under the Train and Test sets.
V OpenFly-Agent
Figure 6: The architecture of OpenFly-Agent. Keyframes at the time of action transitions are selected to extract crucial observations as the history, with corresponding visual tokens compressed to reduce the computational burden.
V-A Dataset Split
Similar to previous works, we divide the dataset into three splits, i.e., Train, Test Seen, Test Unseen. Detailed data distributions are shown in Fig. 5. For the Train split, 7 scenes under the UE rendering engine account for $75.7\%$ of the total 100K data, since UE provides the largest number of scenes, where different amounts of trajectories are sampled according to the scenario area. The 4 scenes created by 3D GS are also the main part of the data, accounting for nearly $20\%$ of the total amount. To ensure visual quality, we only collect data from a high-altitude perspective using Google Earth, which accounts for $4.46\%$. The detailed information of the Test Seen and Test Unseen splits are as follows:
  • Test Seen: 1800 trajectories uniformly sampled from 11 previously seen UE and 3D GS scenarios.
  • Test Unseen: 1200 trajectories uniformly generated from 3 unseen scenarios, i.e., UE-smallcity, 3D GS-sjtu02, and a Los Angeles-like city in GTA V.
V-B Problem Definition
In the aerial VLN task, a UAV is randomly positioned within a 3D environment with its initial pose defined as $P=[x,y,z,\phi,\theta,\psi]$. At each timestamp $t$, the UAV perceives the surrounding environment through an egocentric image as its observation. Guided by natural language instructions, the task involves predicting the next navigation action. Notably, the UAV can utilize either the current observations or the frames from all previous timestamps to make its prediction.
V-C Model Architecture
As shown in Fig. 6, we take OpenVLA [22] as the baseline and design an end-to-end model for aerial VLN. In contrast, our model takes a sequence of images to indicate the observation instead of one image in the original OpenVLA. Moreover, to mitigate visual redundancy between adjacent video frames while maintaining key information, two strategies are proposed, i.e., keyframe selection and visual token merging. First, a series of candidate keyframes are selected. Then, these keyframes are merged temporally before and after the vision encoder, resulting in a compact sequence of visual tokens. Finally, the action decoder discretizes the predicted tokens into uniformly distributed bins, which are subsequently mapped to the 6 action types specific to drones.
V-C1 Keyframe Selection
The length of contextual visual tokens is a major challenge for VLMs when processing videos. Many open-source VLMs use uniform frame sampling [5, 39, 48] to reduce calculation, but this strategy is not suitable for aerial VLN, since it may miss frames containing key landmarks. To address this issue, we adopt a heuristic method to identify keyframes by detecting the change point of the UAV’s movement. We notice that sudden changes in the UAV’s trajectory are often caused by the observation of landmarks, which can serve as cues to determine keyframes. Specifically, we use the movement of the drone over time to draw turning curves, and the frames near the peaks of the wave are selected as candidate keyframes. The resulting data is interpolated and smoothed, forming a wave-like curve that represents the UAV’s movement.
To further ensure the precision of training data, scene segmentation maps collected in Sec. III-A are used on selected frames to detect key landmarks. Frames containing landmarks are selected as keyframes, yielding reasonably accurate results. Note that each sudden change of actions, e.g., from ‘Forward’ to ‘Turn Left’, will produce a set of keyframes. Consequently, we obtain several sets of keyframes for a long trajectory.
V-C2 Visual Token Merging
To further reduce redundant information in keyframes, we design visual token merging, where the core concept is to recognize the similarity between image tokens. It compares adjacent keyframes to merge similar regions and maintains its simplicity by token compression.
TABLE III: Comparison results on the test-seen split.
[…]
For each set of candidate keyframes obtained in the previous selection process, a visual encoder maps each input image to multiple visual tokens, with each token representing the information of an image patch. Considering the potential inter-frame patch redundancy, we take a strategy that similar tokens in subsequent adjacent frames are periodically merged. Specifically, we select the first frame in a keyframe set as the reference, since it usually contains the crucial observation indicating the time for action transition. Then, we densely calculate the cosine similarities between each pair of visual tokens of the reference image and the subsequent image. Next, we merge the tokens with high similarity by averaging them. The unmerged tokens in the subsequent frame will be discarded. The merging operation is iteratively performed until the entire keyframe set has been traversed. Besides, we maintain a memory bank with a capacity of $K$ images, which follows a first-in-first-out (FIFO) policy to retain the latest keyframes.
After the above process, $M$ visual tokens $E=\{e_{1},e_{2},\cdots,e_{M}\}$ are obtained for each set of keyframes. Since aerial VLN requires UAVs to perform long-distance flights based on instructions, we continue to carry out token compress to reduce the computational burden. The compressed visual tokens $E_{c}$ are obtained through grid pooling [27]. Notably, we keep the visual tokens of the current frame uncompressed to capture the latest visual observation, as it contains the most important information for flight action prediction.
V-C3 Action Prediction
[…]
TABLE IV: Comparison results on the test-unseen split.
[…]
TABLE V: Ablation study on the test-seen split. ‘KS’ and ‘VTM’ denote keyframe selection and visual token merging, respectively.
[…]
VI Experiments
VI-A Implementation and Training Details
The proposed OpenFly-Agent is composed of the Dino-SigCLIP (224 $\times$ 224 pixels) as a vision encoder and the pretrained Llama-2 (7B) as a language model. The visual tokens extracted by Dino-SigLIP are aligned to the same dimension as text embeddings using a projection layer, and then fed into the Llama-2 model. The current frame during flight remains 256 tokens and all historical keyframes are compressed into 1 token. The capacity $K$ of the history memory bank is set to 2 in our experiment. For action prediction, the last 256 tokens in the vocabulary are used as special tokens for action representation.
VI-B Evaluation Metrics
Four standard metrics in VLN tasks are adopted to evaluate different methods, i.e., navigation error (NE), success rate (SR), oracle success rate (OSR), and success weighted by path length (SPL). NE measures the average deviation between the UAV’s final stopping point and the ground-truth destination. SR calculates the proportion of successful tasks, where a task is considered successful if the UAV stops within 20 m of the target. In OSR, if any point on the trajectory is within 20 m of the target, the task can be considered successful. SPL calculates the success rate weighted by the ratio of the ground-truth path length to the actually-executed path length.
VI-C Results
VI-C1 Quantitative results
We test the proposed OpenFly-Agent on the test-seen and test-unseen splits and differentiate among easy, moderate, and hard difficulty levels. Multiple VLN methods are compared with ours, where ‘Random’ means an agent randomly selects one action to execute until the ‘stop’ action is chosen, Seq2Seq [23] adopts a simple recurrent model for action sequence prediction, CMA [23] uses a bi-directional LSTM to implement cross-modal attention between instructions and images, AerialVLN [29] combines CMA and a look-ahead guidance strategy to enhance robustness. Navid [56] leverages the strong ability of VLMs and develops a video-based architecture, achieving impressive results in indoor VLN.
Table. III shows quantitative results on the test-seen split. Seq2Seq, CMA, and AerialVLN achieve limited success rates. Counterintuitively, they perform worse on easy samples. We find that their OSR is not low while they are difficult to stop at a short-range distance. In contrast, Navid obtains better results, demonstrating the great potential of VLMs in aerial VLN. Our OpenFly-Agent achieves the best performance for its strong VLM backbone and the proposed strategies, while there is still much room for improvement. Table. IV presents quantitative results on the test-unseen split, illustrating the generalization capabilities of these VLN methods. Similarly, our method achieves better performance, exhibiting a certain degree of robustness. However, all methods are significantly degraded, which indicates models with greater generalization are urgently needed to be developed.
Figure 7: Illustration of an aerial VLN trajectory generated by OpenFly-Agent, which successfully predicts actions following the instruction when encountering landmarks.
Figure 8: Illustration of failure cases. Sometimes our model may misclassify key landmarks or output wrong actions.
VI-C2 Qualitative results
Fig. 7 presents a qualitative result, where our OpenFly-Agent successfully navigates to the destination according to the instruction. It presents a powerful capability in perceiving environments and aligning observations with complex instructions. In addition, Fig. 8 shows two failure cases, where our model fails to identify the landmark or output actions with proper amplitudes.
VI-D Ablation Study
Ablation studies are conducted to evaluate the contribution of the keyframe selection and visual token merging in OpenFly-Agent. Table. V shows the results, where OpenVLA [22] is our baseline. Using only the current frame as an observation makes OpenVLA perform poorly in the aerial VLN task. From ‘History + VTM’ we can see that historical frames significantly improve the success rate. The keyframe selection strategy further increases the SR from 12.7% to 18.5%, demonstrating the effectiveness of key observations. Besides, the comparison between ‘KS’ and ‘KS + VTM’ indicates the great effect of our visual token merging strategy.
VII Limitations and Future Work
The proposed OpenFly platform incorporates various rendering engines/techniques to provide high-quality scenes. Specifically, this is the first attempt to use 3D GS reconstructed scenes to support real-to-sim training and testing, while in the reconstruction of large-scale areas, a few visual artifacts are inevitably present. Future work will focus on exploring more effective reconstruction methods to enhance realism in large-scale scenes. Besides, the proposed OpenFly-Agent is built upon the large VLN model architecture, which is not practical for real-time deployment on UAVs. To address this, future research should focus on developing more efficient architectures and effective quantization techniques.
VIII Conclusion
In this work, we present OpenFly, a platform designed for large-scale data collection in aerial Vision-and-Language Navigation (VLN). OpenFly integrates multiple rendering engines and advanced real-to-sim techniques for data generation, enabling efficient collection of diverse, high-quality aerial VLN data. The resulting large-scale dataset comprises 100k trajectories across 18 distinct scenes, spanning a wide range of altitudes and difficulty levels, which is significantly superior than existing ones. Furthermore, we propose OpenFly-Agent, a keyframe-aware aerial navigation model capable of directly predicting flight actions based on observations and language instructions. Extensive experiments validate the effectiveness of the proposed method, and establishing a comprehensive benchmark for future advancements in aerial navigation.

### v5

@@
-OpenFly: A Versatile Toolchain and Large-scale Benchmark for Aerial Vision-Language Navigation
+OpenFly: A Comprehensive Platform for Aerial Vision-Language Navigation
 Abstract
-Vision-Language Navigation (VLN) aims to guide agents through an environment by leveraging both language instructions and visual cues, playing a pivotal role in embodied AI.
+Vision-Language Navigation (VLN) aims to guide agents by leveraging language instructions and visual cues, playing a pivotal role in embodied AI.
 Indoor VLN has been extensively studied, whereas outdoor aerial VLN remains underexplored.
 The potential reason is that outdoor aerial view encompasses vast areas, making data collection more challenging, which results in a lack of benchmarks.
-To address this problem, we propose OpenFly, a platform comprising a versatile toolchain and large-scale benchmark for aerial VLN.
-Firstly, we develop a highly automated toolchain for data collection, enabling automatic point cloud acquisition, scene semantic segmentation, flight trajectory creation, and instruction generation.
-Secondly, based on the toolchain, we construct a large-scale aerial VLN dataset with 100k trajectories, covering diverse heights and lengths across 18 scenes.
-The corresponding visual data are generated using various rendering engines and advanced techniques, including Unreal Engine, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS).
-All data exhibit high visual quality.
-Particularly, 3D GS supports real-to-sim rendering, further enhancing the realism of the dataset.
-Thirdly, we propose OpenFly-Agent, a keyframe-aware VLN model, which takes language instructions, current observations, and historical keyframes as input, and outputs flight actions directly.
-Extensive analyses and experiments are conducted, showcasing the superiority of our OpenFly platform and OpenFly-Agent.
+To address this problem, we propose OpenFly, a platform comprising various rendering engines, a versatile toolchain, and a large-scale benchmark for aerial VLN.
+Firstly, we integrate diverse rendering engines and advanced techniques for environment simulation, including Unreal Engine, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS).
+Particularly, 3D GS supports real-to-sim rendering, further enhancing the realism of our environments.
+Secondly, we develop a highly automated toolchain for aerial VLN data collection, streamlining point cloud acquisition, scene semantic segmentation, flight trajectory creation, and instruction generation.
+Thirdly, based on the toolchain, we construct a large-scale aerial VLN dataset with 100k trajectories, covering diverse heights and lengths across 18 scenes.
+Moreover, we propose OpenFly-Agent, a keyframe-aware VLN model emphasizing key observations during flight.
+For benchmarking, extensive experiments and analyses are conducted, evaluating several recent VLN methods and showcasing the superiority of our OpenFly platform and agent.
 The toolchain, dataset, and codes will be open-sourced.
 Figure 1: Overview of OpenFly.
-This work consists of an automatic toolchain for data generation, a large-scale aerial VLN dataset comprising 100K trajectories and instructions, and a keyframe-aware VLN model emphasizing key observations.
-I Introduction
-Figure 2: Framework of the automatic data generation platform.
-Various rendering engines and simulators are first integrated, providing diverse high-quality scenes.
+This work consists of the integration of 4 rendering engines, an automatic toolchain for data generation, a large-scale aerial VLN dataset comprising 100K trajectories and instructions, and a keyframe-aware VLN model emphasizing key observations.
+Keywords: Aerial Vision-Language Navigation (VLN), benchmark, toolchain
+1 Introduction
+Figure 2: Framework of the automatic data generation.
+Various rendering engines are first integrated, providing diverse high-quality scenes.
 Built on these, several interfaces and tools are developed, enabling automated generation of trajectories and instructions.
@@
 Current datasets for aerial VLN remain relatively small, with only around 10k trajectories, which is far behind embodied manipulation datasets.
-For example, Open X-Embodiment [36] has collected 1M episodes of manipulation data, which has significantly accelerated the development of vision-language-action (VLA) models [38, 54, 35, 55].
-To address these issues, we propose OpenFly, a comprehensive platform consisting of a versatile toolchain and a large-scale benchmark for the aerial VLN task.
-To enhance data diversity, the platform is established on various widely-used rendering engines and advanced techniques, i.e., UE, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS).
-This enables us to utilize a wide range of assets as shown in Fig.
+For example, Open X-Embodiment [19] has collected 1M episodes of manipulation data, which has significantly accelerated the development of vision-language-action (VLA) models.
+To address these issues, we propose OpenFly, a comprehensive platform consisting of diverse rendering engines, a versatile toolchain, and a large-scale benchmark for the aerial VLN task.
+To enhance data diversity, the platform is established on various widely-used rendering engines and advanced techniques, i.e., UE, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS), enabling us to utilize a wide range of assets as shown in Fig.
 1.
-In particular, we use UAVs for automated patrols to capture numerous real-world images and integrate 3D GS technology into our platform, through which we reconstruct realistic 3D scenes.
-This empowers our platform to support a real-to-sim method, enhancing the realism of data.
-To improve the efficiency of data collection, we develop a versatile toolchain for automated aerial VLN data generation, as depicted in Fig.
+In particular, we use UAVs to capture numerous real-world images and integrate 3D GS technology into our platform to reconstruct realistic 3D scenes, empowering real-to-sim simulation.
+To improve the efficiency of data collection, we develop a versatile toolchain for automated aerial VLN data generation as depicted in Fig.
 2.
@@
 Benefitting from our toolchain, we are able to quickly construct a dataset of 100k samples, significantly larger than existing datasets.
-Besides, we propose OpenFly-Agent, a keyframe-aware aerial VLN model trained on the OpenFly dataset.
-It is developed based on a VLA model OpenVLA [22] and incorporates an adaptive frame-level token-sampling mechanism to mitigate visual distraction caused by redundant historical observations.
-Note that this sampling mechanism is crucial for the visual encoding of aerial VLN, as UAVs fly quickly and observations change rapidly.
+Besides, we propose OpenFly-Agent, a keyframe-aware aerial VLN model incorporating an adaptive frame-level sampling mechanism to emphasize critical observations.
+Note that this keyframe sampling mechanism is crucial for the visual encoding of aerial VLN, as UAVs fly quickly and observations change rapidly.
 Extensive experiments are conducted on the OpenFly dataset to evaluate numerous methods, establishing a comprehensive benchmark for the aerial VLN tasks.
@@
   • We build OpenFly on multiple rendering engines and develop a versatile toolchain, enabling the automatic generation of data with both great diversity and efficiency.
-  • We have constructed a novel aerial VLN benchmark comprising 100k trajectories across 18 high-quality scenes. To the best of our knowledge, this is the largest aerial VLN benchmark to date.
+  • We have constructed a novel aerial VLN benchmark comprising 100k trajectories across 18 high-quality scenes.
+To the best of our knowledge, this is the largest aerial VLN benchmark to date, and users can collect more customized data using the OpenFly platform.
   • We propose OpenFly-Agent, a keyframe-aware VLN model. Extensive experiments demonstrate its superior performance compared to other methods.
-II Related Works
-II-A Simulators for Embodied AI
-In recent years, many simulators related to embodied AI have been developed, e.g., MuJoCo [46], Habitat [42], Pybullet [11], Matterport3D [6], OpenAI Gym [4], and Issac Gym [32].
-They allow researchers to learn and validate algorithms in safe virtual environments, significantly advancing the development of embodied intelligence.
-However, most of them are designed for indoor robotic manipulation and navigation, making them unsuitable as simulators for aerial VLN tasks.
-Gazebo and AirSim [44] are two popular UAV simulators, while Gazebo suffers from low rendering quality.
-Combining AirSim and UE is a potential solution to construct aerial VLN scenes.
-Unfortunately, the official maintenance and updates for AirSim have been discontinued, resulting in a limited availability of compatible assets.
-II-B Vision-Language Navigation Datasets
+2 Related Works
+2.1 Vision-Language Navigation Datasets
 Numerous datasets have been proposed to accelerate the VLN task.
 R2R [3] focuses on evaluating agents in unseen buildings and provides discrete navigation options.
-RxR [25] is a densely annotated VLN dataset, supporting multilingual instructions.
+RxR [4] provides a more densely annotated VLN dataset.
 TouchDown [1] and REVERIE [2] have each contributed a dataset from real-life environments, which requires a ground-based agent to follow instructions and find a target object.
-CVDN [45] presents a cooperative VLN dataset where agents can access the dialog history of human cooperation for inference.
-All the above datasets are graph-based where navigable points are predefined.
+CVDN [5] presents a cooperative VLN dataset where agents can access the history of human cooperation for inference.
+All the above datasets are graph-based, where navigable points are predefined.
 LANI [7] and VLN-CE [6] propose the VLN task in continuous outdoor/indoor environments, enabling agents to move freely to any unobstructed point.
@@
 AerialVLN [17] and OpenUAV [18] integrate AirSim and UE to create VLN scenes where pilots can control UAVs to generate various trajectories.
-II-C Vision-Language Navigation Methods
+2.2 Vision-Language Navigation Methods
 VLN methods require agents to understand and complete language instructions according to visual observations.
@@
 In contrast,  [31, 32, 33, 16] explore VLN in continuous environments for more realistic applications.
-More recently, a few works attempt to address the problem of aerial VLN based on UAVs, among which AerialVLN [29] proposes a lookahead guidance method to generate more reasonable ground-truth trajectories during training, and STMR [14] designs a matrix representation to improve the spatial reasoning of LLMs.
-III OpenFly Data Generation Platform
-Figure 3: High-quality examples from different rendering engines and techniques, including several large cities such as Shanghai, Guangzhou, Los Angeles, Osaka, and etc., cover an area of over a hundred square kilometers in total.
-3D GS provides five large campus scenes, further enhancing the diversity and realism of the data.
-In this section, we first describe several basic simulators and data resources, and then present the developed toolchain.
-The framework of the whole automatic data generation platform is illustrated in Fig.
-2.
-III-A Basic Simulators and Data Resources
-To collect a wide range of high-quality and realistic simulation data, we source the dataset from multiple rendering engines integrated with various simulators.
+More recently, a few works attempt to address the problem of aerial VLN based on UAVs, among which AerialVLN [17] proposes a lookahead guidance method to generate more reasonable ground-truth trajectories during training.
+STMR [34] designs a matrix representation to improve the spatial reasoning of LLMs.
+OpenUAV [18] incorporates extra human assistance aligned with the ground truth (GT) trajectory to guide the navigation.
+3 Automatic Data Generation
+In this section, we first describe several rendering engines and data resources, then present the developed toolchain. The whole automatic data generation framework is illustrated in Fig. 2.
+3.1 Rendering Engines and Data Resources
+Unreal Engine.
+UE is a rendering engine capable of providing highly realistic virtual environments.
+We meticulously select 8 urban scenes, covering more than $100km^{2}$ areas.
+These scenes include a variety of assets such as buildings, streets, traffic lights, vehicles, and pedestrians, encompassing a wide range of architectural styles.
+GTA V.
+GTA V is an open-world game that is frequently used by researchers due to its highly realistic environment.
+The game features a meticulously crafted cityscape modeled after Los Angeles, encompassing various buildings and locations such as skyscrapers, gas stations, parks, and plazas.
+Google Earth.
+Google Earth is a virtual globe software, which builds a 3D earth model by integrating satellite imagery and aerial photographs.
+From this engine, we select four urban scenes covering a total area of $53.60km^{2}$, i.e., Berkeley, Osaka, Washington, D.C., and St.
+Louis.
+3D Gaussian Splatting.
+3D GS excels in realistic scene reconstruction by capturing intricate textures.
+It enables the creation of highly detailed, lifelike virtual environments, significantly improving the realism of digital scenes.
+Due to these features, we use a hierarchical 3D GS method [35] to reconstruct and render multiple real scenes.
+Practically, we gathered data from five campuses across three universities using UAVs.
+These campus scenes include various types and styles of landmarks, such as libraries, bell towers, waterways, lakes, playgrounds, construction sites, and lawns.
+More details of rendering engines, data resources, and examples can be found in Appendix A.
+3.2 Toolchain for Automatic Data Collection
+To achieve automatic data generation, we first integrate the above rendering engines and design three unified interfaces to control the agent movement and acquire sensor data (presented in Appendix C).
+Based on these interfaces, we further develop a toolchain, streamlining point cloud acquisition, scene semantic segmentation, trajectory creation, and instruction generation.
+3D Point Cloud Acquisition.
+Our OpenFly provides two methods to reconstruct the point cloud map for different scenes.
+1) Rasterized Sampling Reconstruction: For UE and GTA V scenes, we customize rasterized sampling points at appropriate resolutions, and then use the developed interface to obtain the local point cloud at the sampling points and stitch them for the entire scene.
+2) Image-Based Sparse Reconstruction: In 3D GS, the scene reconstruction process begins with the open-source COLMAP [36] framework, which generates a sparse point cloud from input images.
+We directly export and use the point clouds obtained from this step.
+Scene Semantic Segmentation.
+VLN requires meaningful landmarks as navigation targets.
+Thus, we offer three semantic segmentation methods to identify landmarks.
+1) 3D Scene Understanding: A sequence of top-down views of the scene is captured in a rasterized format, followed by the off-the-shelf Octree-Graph [37] to extract semantic 3D instances.
+2) Point Cloud Projection and Contour Extraction: We acquire the point cloud of a scene and project the voxelized point cloud onto the ground.
+For each instance, its contour is segmented using OpenCV, and the maximum height of its points is used as the final height.
+Additionally, semantic annotations are obtained by feeding the segmented instances to GPT-4o for caption.
+3) Manual Annotation: When the point cloud quality of a scene is low or finer segmentation is required, OpenFly provides an interface for manually annotating instances and semantics within the point cloud.
+Users can choose these methods flexibly based on their requirements.
+The corresponding results are shown in our appendix.
+Automatic Trajectory Generation.
+Leveraging the point cloud map and segmentation tools, OpenFly can automatically generate VLN trajectories using the following method.
+First, a global voxel map $M_{global}$ is constructed from the scene point cloud.
+Second, a landmark is randomly chosen as the target.
+A starting point is selected at a certain distance from the landmark, and a point close to the landmark is chosen as the endpoint.
+Third, A collision-free trajectory is generated using the A* [38] pathfinding algorithm based on $M_{global}$ and customized action space.
+By repeatedly selecting the endpoint as the new starting point, complex trajectories can be generated.
+Finally, utilizing OpenFly’s interface, images corresponding to the trajectory points are obtained as visual observations.
+A more comprehensive introduction has been included in Appendix E.
+Automatic Instruction Generation.
+Previous research has predominantly relied on manual annotation to generate trajectory instructions, which is not only costly but also limits the scalability of datasets.
+To address this issue, we propose a highly automated instruction generation method based on VLMs, e.g., GPT-4o.
+A straightforward method would be to submit all images to VLMs to analyze the trajectory and generate instructions.
+However, using all images introduces considerable computational overhead and causes significant difficulties for a VLM to understand the trajectory.
+Additionally, we find the ‘Forward’ action usually occupies a larger proportion of a flight trajectory, with ‘Turn Left/Turn Right’ or ‘Ascend/Descend’ actions taken when encountering key landmarks.
+Therefore, we split the complete trajectory into multiple sub-trajectories according to the occurrence of non-consistent actions, extracting key actions and images for processing.
+Notably, slight angle adjustments often occur during flight to change the direction subtly, and a slight ‘Turn’ will be merged with subsequent ‘Forward’ actions.
+We submit the action sequence and the last captured image of each sub-trajectory to a VLM to generate a sub-instruction of both the action and the landmark.
+All sub-instructions of the same trajectory are then processed by an LLM to integrate into a complete instruction.
+The proposed strategy significantly improves the instruction accuracy than directly inputting all trajectory images to a VLM.
+To further verify the data quality, we randomly select 3K samples from the entire dataset according to the data distribution in Sec.
+4.2.
+After manually inspecting these samples, we find that they reach a high qualification rate of 91%.
+The problematic data involves some vague descriptions, but it is still considered acceptable by examiners.
+More details of instruction generation and data quality control are provided in Appendix F and G.
+4 Dataset Analysis
+Figure 3: Statistical analysis of the generated data. (a) Length and height distributions of trajectories. (b) Action distributions. (c) Word cloud of nouns. (d) Word cloud of verbs.
+4.1 Trajectory and Instruction Analysis
+Using our toolchain, we collect a dataset of 100K trajectories, which is much larger than other VLN datasets.
+Compared with ground-based VLN, the aerial VLN task has more motion dimensions.
+Therefore, we set different trajectory lengths and flight heights to obtain rich data.
 Fig.
-3 showcases several examples obtained from these rendering engines/techniques.
-Unreal Engine + AirSim/UnrealCV.
+3(a) exhibits the distribution of these data, with their lengths ranging from 0 to 300 meters, and the heights ranging from 0 to 150 meters.
+Notably, we follow the mainstream methods [6, 17] to use discrete actions, e.g., ‘Forward’ and ‘Turn left’, for trajectory generation, where the step size is 3m for ‘Forward’ action, making its proportion much higher than that of ‘Turn’ and altitude adjustments.
+This highly unbalanced action distribution might cause the VLN model to overfit the dominant action.
+To alleviate this problem, we employ three granularities of ‘Forward’ action during trajectory generation, i.e., 3m, 6m, and 9m.
+Fig.
+3(b) presents the action distribution of our dataset.
+It should be noted that the OpenFly platform also supports trajectory generation with continuous waypoints.
+For instruction analysis, the vocabulary size of our dataset is 15.6K, and the average length of instructions is 59.
+Fig.
+3(c) and Fig.
+3(d) illustrate the word clouds of nouns and verbs, where ‘building’, ‘windows’, and ‘skyscraper’ are the most common references, and ‘proceed’ and ‘turn’ are the mostly used verbs for VLN.
+Due to the space limitation, we put more details of our dataset and other ones as well as more instruction analyses in Appendix H.
+4.2 Dataset Split
+Similar to previous works, we divide the dataset into three splits, i.e., Train, Test Seen, Test Unseen.
+For the Train split, 7 scenes under the UE rendering engine account for $75.7\%$ of all data, since UE provides the largest number of scenes, where different amounts of trajectories are sampled according to the areas of scenes.
+The 4 scenes created by 3D GS are also the main part of the data, accounting for nearly $20\%$ of the total amount.
+To ensure visual quality, we only collect data from a high-altitude perspective using Google Earth, which accounts for $4.46\%$.
+The Test Seen data consists of 1800 trajectories uniformly sampled from 11 previously seen UE and 3D GS scenes, and the Test Unseen data comprises 1200 trajectories uniformly generated from 3 unseen scenes, i.e., UE-smallcity, 3D GS-sjtu02, and a Los Angeles-like city in GTA V.
+Detailed data distributions are shown in Appendix H.
+5 OpenFly-Agent
+Figure 4: The architecture of OpenFly-Agent.
+Keyframes at the time of action transitions are selected to extract crucial observations as the history, with corresponding visual tokens compressed to reduce the computational burden.
+In this section, we present OpenFly-Agent, an aerial VLN model that builds upon the OpenVLA [39] baseline, since OpenVLA and aerial VLN share the same pipeline, i.e., taking images and instructions as input and generating actions.
+Moreover, OpenVLA is trained on 1M data, having strong abilities in instruction-following and reasoning, which is critical for VLN.
+In contrast, our model takes a sequence of images to indicate the observation history instead of one image in the original OpenVLA.
+Additionally, to mitigate visual redundancy between adjacent video frames while maintaining key information, two strategies are proposed, i.e., keyframe selection and visual token merging.
+First, a series of candidate keyframes is selected.
+Then, these keyframes are merged temporally before and after the vision encoder, resulting in a compact sequence of visual tokens.
+Finally, the action decoder discretizes the predicted tokens to 6 action types specific to UAVs.
+Fig.
+4 illustrates the architecture of our OpenFly-Agent, with the details of two strategies described below.
+5.1 Keyframe Selection
+The length of contextual visual tokens is a major challenge for VLMs when processing videos.
+Many open-source VLMs use uniform frame sampling [40, 41, 42] to reduce calculation, but this strategy is not suitable for aerial VLN, since it may miss frames containing key landmarks.
+We notice that sudden changes in the UAV’s trajectory are often caused by the observation of landmarks in both our dataset and AerialVLN [17], which can serve as cues to determine keyframes.
+Therefore, we adopt a heuristic method to select keyframe candidates near the change point of the UAV’s movement.
+To further ensure the precision of training data, scene segmentation maps collected in Sec.
+3.2 are used on selected frames to detect landmarks.
+Frames containing landmarks are selected as keyframes, yielding reasonably accurate results.
+Note that each sudden change of actions, e.g., from ‘Forward’ to ‘Turn Left’, will produce a set of keyframes.
+Consequently, we obtain several sets of keyframes for a long trajectory.
+5.2 Visual Token Merging
+To further reduce redundant information in keyframes, we design visual token merging, where the core concept is to recognize the similarity between image tokens.
+It compares adjacent keyframes to merge similar regions and maintains its simplicity by token compression.
+Table 1: Comparison results on the test set. ‘Random’ means randomly selecting one action to execute until the ‘stop’ action is chosen. All models are retrained using our dataset.
+Method test-seen test-unseen
+[…]
+Random 242m 0.7% 0.8% 0% 301m 0.1% 0.1% 0%
+Seq2Seq[6] 291m 0.1% 3.2% 0.6% 291m 0.1% 0.9% 0.6%
+CMA[6] 144m 5.7% 30.0% 5.1% 268m 0.7% 6.6% 0.7%
+AerialVLN[17] 130m 6.6% 40.8% 7.0% 227m 2.3% 19.8% 2.2%
+Navid[16] 142m 9.9% 24.3% 3.9% 231m 4.2% 27.0% 3.5%
+NaVila[43] 140m 14.2% 41.7% 6.7% 226m 4.5% 25.1% 2.9%
+OpenFly-Agent (Ours) 115m 18.5% 50.9% 12.2% 198m 5.1% 27.3% 3.5%
+For each set of candidate keyframes obtained in the previous selection process, a visual encoder maps each input image to multiple visual tokens, with each token representing the information of an image patch.
+Considering the potential inter-frame patch redundancy, we take a strategy that similar tokens in subsequent adjacent frames are periodically merged.
+Specifically, we select the first frame in a keyframe set as the reference, since it usually contains the crucial observation indicating the time for action transition.
+Then, we densely calculate the cosine similarities between each pair of visual tokens of the reference image and the subsequent image.
+Next, we merge the tokens with high similarity by averaging them.
+The unmerged tokens in the subsequent frame will be discarded.
+The merging operation is iteratively performed until the entire keyframe set has been traversed.
+Besides, we maintain a memory bank with a capacity of $K$ images, which follows a first-in-first-out (FIFO) policy to retain the latest keyframes.
+Since aerial VLN requires UAVs to perform long-distance flights based on instructions, we continue to conduct token compression within each keyframe to reduce the computational burden.
+The compressed visual tokens are obtained through grid pooling [44].
+Notably, we keep the visual tokens of the current frame uncompressed to capture the latest visual observation, as it contains the most important information for action prediction.
+6 Experiments
+In this section, we first introduce evaluation metrics, and then present the experimental results and ablation studies. The implementation details of the OpenFly-Agent can be found in Appendix I.
+6.1 Evaluation Metrics
+Four standard metrics in VLN tasks are adopted to evaluate different methods, i.e., navigation error (NE), success rate (SR), oracle success rate (OSR), and success weighted by path length (SPL).
+NE measures the average deviation between the UAV’s final stopping point and the ground-truth destination.
+SR calculates the proportion of successful tasks, where a task is considered successful if the UAV stops within 20 m of the target [17].
+In OSR, if any point on the trajectory is within 20 m of the target, the task can be considered successful.
+SPL calculates the success rate weighted by the ratio of the ground-truth path length to the actually-executed path length.
+Table 2: Ablation study on the test-seen split. ‘KS’ and ‘VTM’ denote keyframe selection and visual token merging, respectively.
+[…]
+OpenVLA (baseline) 196m 3.6% 12.4% 1.9%
+History + VTM 215m 12.7% 39.6% 6.4%
+KS 275m 6.2% 24.1% 5.6%
+KS + VTM 115m 18.5% 50.9% 12.2%
+6.2 Quantitative Results
+We evaluate the proposed OpenFly-Agent and multiple VLN methods on the test set, with quantitative results listed in Table 1, where Seq2Seq, CMA, and AerialVLN achieve limited success rates.
+In contrast, Navid [16] and NaVila [43] are two most recent VLN methods, obtaining better results and demonstrating the great potential of VLMs in aerial VLN.
+Our OpenFly-Agent achieves the best performance for its strong VLM backbone and the proposed strategies, while aerial VLN is an emerging and challenging task, and there is still much room for improvement.
+The results on the test-unseen split indicate the generalization abilities of these methods.
+Similarly, our method achieves better performance, exhibiting a certain degree of robustness.
+However, all methods are significantly degraded, indicating that more powerful models are urgently needed to be developed.
+More qualitative results and real-world experiments are shown in Appendix J and K.
+6.3 Ablation Study
+Ablation studies are conducted to evaluate the contribution of the keyframe selection and visual token merging in OpenFly-Agent.
+Table.
+2 shows the results, where OpenVLA [39] is our baseline.
+Using only the current frame as an observation makes OpenVLA perform poorly in the aerial VLN task.
+From ‘History + VTM’ we can see that historical frames significantly improve the success rate.
+The keyframe selection strategy further increases the SR from 12.7% to 18.5%, demonstrating the effectiveness of key observations.
+Besides, the comparison between ‘KS’ and ‘KS + VTM’ indicates the great effect of our visual token merging strategy.
+7 Limitations and Future Work
+The proposed OpenFly platform incorporates various rendering engines/techniques to provide high-quality scenes.
+Specifically, this is the first attempt to use 3D GS reconstructed scenes to support real-to-sim training and testing, while in the reconstruction of large-scale areas, a few visual artifacts are inevitably present.
+Future work will focus on exploring more effective reconstruction methods to enhance realism in large-scale scenes.
+Besides, the proposed OpenFly-Agent is built upon the large VLM model architecture, which is not practical for real-time deployment on UAVs.
+To address this, future research should focus on developing more efficient architectures and quantization techniques.
+8 Conclusion
+In this work, we present OpenFly, a platform designed for large-scale data collection in aerial Vision-and-Language Navigation (VLN).
+OpenFly integrates multiple rendering engines and provides an automatic toolchain for data generation, enabling efficient collection of diverse, high-quality aerial VLN data.
+The resulting large-scale dataset comprises 100k trajectories across 18 distinct scenes, spanning a wide range of altitudes and lengths, which is significantly superior than existing ones.
+Furthermore, we propose OpenFly-Agent, a keyframe-aware aerial navigation model capable of directly predicting flight actions based on observations and language instructions.
+Extensive experiments validate the effectiveness of the proposed method, and establish a comprehensive benchmark for future advancements in aerial VLN.
+Appendix
+Appendix A More Details of Rendering Engines and Data Resources
+In this section, more information about the used rendering engines and data resources are detailed, with several high-quality examples illustrated in Fig. 5.
+Unreal Engine.
 UE is a rendering engine capable of providing highly realistic interactive virtual environments.
@@
 To realize a highly efficient data collection in simulated scenes, we modify the UE5 project to a C++ project, integrate the UnrealCV plugin, and package executables for multiple systems like Windows and Linux.
-GTA V + Script Hook V.
+GTA V.
 GTA V is an open-world game that is frequently used by computer vision researchers due to its highly realistic and dynamic virtual environment.
@@
 From the interface, we can control a virtual agent to collect the required data in an arbitrary pose and angle in the game.
-Google Earth + Google Earth Studio.
+Google Earth.
 Google Earth is a virtual globe software, which builds a 3D earth model by integrating satellite imagery, aerial photographs, and Geographic Information System (GIS) data.
@@
 In order to efficiently plan the route, we develop a function that automatically draws the flight trajectory in Google Earth Studio according to the selected area and predefined photo interval.
-3D Gaussian Splatting + SIBR viewers.
+3D Gaussian Splatting.
 As a highly realistic reconstruction method, hierarchical 3D GS  [35] employs a hierarchical training and display architecture, making it particularly suitable for rendering large-scale areas.
@@
 These campus scenes include various types and styles of landmarks, such as libraries, bell towers, waterways, lakes, playgrounds, construction sites, and lawns.
-The detailed information for the five campuses is presented in Table I.
-More details of the 3D GS data collection can be found in our supplementary material.
+The detailed information for the five campuses is presented in Table 3.
 SIBR [46] viewers is a rendering tool designed for the 3D GS project, enabling visualization of a scene from arbitrary viewpoints.
@@
 Building upon SIBR viewers, we developed an HTTP RESTful API that generates RGB images from arbitrary poses, simulating a UAV’s perspective.
-TABLE I: Different 3D GS Scenes
+Figure 5: High-quality examples from different rendering engines and techniques, including several large cities such as Shanghai, Guangzhou, Los Angeles, Osaka, and etc., cover an area of over a hundred square kilometers in total.
+3D GS provides five large campus scenes, further enhancing the diversity and realism of the data.
+Table 3: Different 3D GS Scenes
 Campus Name Images Area
@@
 SJTU (Minghang-West Zone) 9536 $0.95km^{2}$
-III-B Toolchain for Automatic Data Colleciton
-To achieve automatic data generation, we integrate the above five simulators and design three unified interfaces, i.e., the agent movement interface, the lidar data acquisition interface, and the image acquisition interface, allowing an agent to interact with any scene.
-Based on these interfaces, we further develop a toolchain, including 3D point cloud acquisition, scene semantic segmentation, automatic trajectory generation, and instruction generation.
-The framework of the whole data generation platform is illustrated in Fig.
-2, with details of these interfaces and tools elaborated below.
-Unified Interfaces.
-1) Agent Movement Interface: We design a CoorTrans module, which implements a customized pose transformation matrix and scaling function to unify all simulator coordinate systems into a meter-based FLU (Front-Left-Up) convention.
-This interface enables precise agent positioning among regular scenes, point clouds, and scene segmentations, ensuring consistency and facilitating automatic trajectory generation.
-2) Lidar Data Acquisition: For different simulators, point cloud data is acquired through different methods, including lidar sensor collection, depth map back-projection, and image feature matching.
+Appendix B Details of 3D GS Data Collection
+From the UAV’s perspective, choosing the appropriate shooting altitude poses a dilemma, i.e., if the altitude is too low, the sparse point cloud generated during the initialization of the 3D GS reconstruction will be suboptimal, due to insufficient feature point matches between photos.
+In contrast, if the altitude is too high, the Gaussian reconstruction will result in an overly coarse training of details.
+After multiple attempts, the data collection plan using the M30T was determined as follows.
+For large-scale block scenes, oblique photography is performed at approximately twice the average building height using the default parameters of the M30T’s wide-angle camera, with a tilt angle of -45°.
+For landmark buildings with heights significantly different from the average height, additional targeted data collection is conducted at twice their height.
+This altitude setting can, to a certain extent, ensure both higher-quality point cloud initialization and Gaussian splatting training.
+Appendix C Unified Interfaces
+For data collection, we integrate all rendering engines and design three unified interfaces, i.e., the agent movement interface, the lidar data acquisition interface, and the image acquisition interface, allowing an agent to move and perceive the environment within any scene.
+  • Agent Movement Interface: We design a CoorTrans module, which implements a customized pose transformation matrix and scaling function to unify all coordinate systems into a meter-based FLU (Front-Left-Up) convention.
+This interface enables precise agent positioning, ensuring consistency and facilitating automatic trajectory generation.
+  • Lidar Data Acquisition Interface: Lidar data is crucial for scene occupancy perception and essential for trajectory generation.
+Our platform supports different lidar data acquisition methods, including lidar sensor collection, depth map back-projection, and image feature matching.
 We develop a unified interface to integrate these methods and leverage the proposed CoorTrans module to align all data to the same FLU coordinate system.
-3) Image Acquisition Interface: We integrate HTTP RESTful and TCP/IP protocols to form a unified image request interface, allowing image data to be obtained from any location with flexible resolutions and agent viewpoints.
-3D Point Cloud Acquisition.
-For different simulators, we provide two methods to reconstruct the point cloud map of an entire scene.
-1) Rasterized Sampling Reconstruction: For the UE5 + UnrealCV simulator, the MatrixCity [28] project offers a convenient rasterized sampling solution.
-We use the aforementioned lidar data acquisition interface to obtain the local point cloud at the sampling points.
-Since these data are already aligned within the same coordinate system, the point cloud map of the entire scene can be constructed by simply stitching local point clouds.
-For the UE4 + AirSim and GTA V simulators, we customize rasterized sampling points at appropriate resolutions, and perform sampling and reconstruction using the agent movement and lidar data acquisition interfaces.
-2) Image-based Sparse Reconstruction: In 3D GS, the scene reconstruction process begins with the open-source COLMAP [43] framework, which geneoverlearates a sparse point cloud from input images.
-We directly export and use the point clouds obtained from this step.
+  • Image Acquisition Interface: We integrate HTTP RESTful and TCP/IP protocols to form a unified image request interface, allowing image data to be obtained from any location with flexible resolutions and viewpoints.
+Figure 6: Results of our point cloud acquisition and semantic segmentation.
+(a) Rasterized sampling point cloud reconstruction.
+(b) Image-based point cloud reconstruction.
+(c) Semantic 3D scene segmentation.
+(d) Point cloud projection and contour extraction.
+Appendix D Results of Point Cloud Acquisition and Semantic Segmentation
+This section presents results regarding the point cloud acquisition and scene semantic segmentation, as shown in Fig. 6.
+Point Cloud Acquisition.
+OpenFly integrates data resources from different rendering engines.
+In order to obtain point cloud information from different scenes, we provide two point cloud acquisition methods.
+1) For UE and GTAV scenes, we provide a tool that utilizes rasterized point cloud sampling and reconstruction to obtain a global point cloud.
+The results can be seen in Fig.
+6(a).
+2) For 3D GS scenes, we use COLMAP [36] to obtain relatively sparse point data from images, as shown in Fig.
+6(b).
+Although the point cloud generated by this method is relatively sparse, it provides sufficient coverage information.
 Scene Semantic Segmentation.
-Vision-and-Language Navigation (VLN) requires meaningful landmarks as navigation targets.
-We perform semantic segmentation on four types of simulation scenes using the following three methods.
-1) 3D Scene Understanding: A sequence of top-down views of the scene is captured in a rasterized format.
-We then use Octree graph [50] to extract 2D mask proposals, which are subsequently projected into the 3D point cloud space to generate semantic 3D segments.
-2) Point Cloud Projection and Contour Extraction: We first acquire the point cloud of a scene, then project the voxelized point cloud onto a projection plane slightly above the ground.
-Using OpenCV, we perform a series of operations on the projected image, including binarization, erosion, and contour extraction, to obtain multiple instances along with their 2D coordinates.
-For each instance, the maximum height of the points within its neighborhood is used as the final height.
-This method provides a more computationally efficient yet coarser segmentation compared to the first approach, allowing users to choose based on their requirements.
-3) Manual Annotation: When the point cloud quality of a scene is low or finer segmentation is required, we provide a method for annotating instances in the point cloud space by mouse clicks, based on ROS2 and RVIZ2.
-Users can annotate instances directly in the point cloud space using mouse clicks to define landmarks of interest for the task.
-This method is applicable to four simulators, i.e., UE + UnrealCV, UE + Airsim, GTA V, and 3D GS.
-Automatic Trajectory Generation.
-Leveraging the aforementioned point cloud map and segmentation tools, OpenFly can generate trajectories using the following two methods.
-1) Path search based on customized action space: First, a global hash voxel map $M_{global}$ and a bird’s eye view (BEV) occupancy map $M_{bev}$ are constructed from the scene point cloud $P$.
+To meet the requirements of different data resources and different segmentation granularities, our OpenFly offers three methods for obtaining 3D semantic segmentation of scenes.
+Here, we present results of segmentation methods other than manual annotation.
+1) Fig.
+6(c) illustrates the semantic segmentation results based on the off-the-shelf 3D scene understanding method Octree-Graph [37].
+This method provides more granular results.
+2) The result of semantic segmentation via point cloud projection and contour extraction is shown in Fig.
+6(d).
+This method leverages high-precision point clouds to achieve instance segmentation for structures like buildings and trees, which directly contact the ground.
+Appendix E A Comprehensive Introduction to Automatic Trajectory Generation
+Leveraging the point cloud map and segmentation tools, OpenFly offers two methods for trajectory generation for different scenes.
+1) Path search based on customized action space: First, a global voxel map $M_{global}$ and a bird’s eye view (BEV) occupancy map $M_{bev}$ are constructed from the scene point cloud.
 Second, the flight altitude is randomly selected within the user-defined height range, and landmarks that are not lower than the height threshold $H_{\tau}$ are chosen as targets.
@@
 2) Path search based on grid: Google Map data does not allow image retrieval at arbitrary poses in the space.
-Thus we rasterize a pre-selected area and collect images from each grid point in all possible orientations.
+Thus, we rasterize a pre-selected area and collect images from each grid point in all possible orientations.
 Starting and ending points are chosen within the grid points to generate trajectories.
 Corresponding images for these trajectory points are then selected from the pre-collected image set.
-Automatic Instruction Generation.
-Previous research has predominantly relied on manual annotation to generate trajectory instruction, which is not only costly but also limits the scalability of datasets.
-To address this issue, we propose a highly automated language instruction generation method based on VLMs, e.g., GPT-4o.
-A straightforward method would be to submit all images to VLMs to analyze the trajectory and generate instructions.
-However, using all images introduces significant computational overhead and leads to information redundancy.
-For example, the ‘Forward’ action usually occupies a larger proportion of a flight trajectory, with ‘Turn Left/Turn Right’ or ‘Ascend/Descend’ actions taken when encountering key landmarks.
-Based on the above findings, we split the complete trajectory into multiple sub-trajectories based on the occurrence of non-consistent actions, extracting key actions and images for processing and subsequent integration.
-Notably, slight angle adjustments often occur during flight to change the direction subtly, and a ‘slightly Turn Left/Right’ will be merged with subsequent ‘Forward’ actions.
-Specifically, suppose that the trajectory action sequence is [1, 1, 2, 1, 1, 2, 2, 1, 0], where 1, 2, and 0 denote ‘Forward’, ‘Turn Left’, and ‘Stop’, respectively.
-This trajectory would be split into four sub-trajectories, i.e., [1, 1], [2, 1, 1], [2, 2], and [1, 0].
-The second sub-trajectory involves ‘slightly turn left’ and ‘move forward’.
-We submit the action sequence and the last captured image of each sub-trajectory to a VLM to generate descriptions of both action and landmarks.
-The sub-instructions are obtained similar to the following format:
-  • ‘Move forward’ to ‘Landmark 1’.
-  • ‘Slightly turn left and move forward’ to ‘Landmark 2’.
-  • ‘Turn left’ towards ‘Landmark 3’.
-  • ‘Move forward’ to ‘Landmark 4’.
-These sub-instructions are then processed by a VLM/LLM again, where they are integrated into coherent and complete instructions.
-The detailed prompt used for the VLM, along with the complete responses, is provided in the supplementary material.
-III-C Quality Control.
+Appendix F Details of Instruction Generation
+Except for the instruction generation described in the main paper.
+The remaining process is mainly divided into two parts: landmark feature extraction and sub-instruction fusion.
+A simplified prompt to the VLM and the corresponding response are probably like this.
+  • Get Landmark features.
+    System Prompt: You are an assistant who is proficient in image recognition.
+You can accurately identify the object in the picture and its characteristics that are different from the surrounding objects.
+I will give you the three final images you will see.
+Please focus on the last image and tell me the features of the target building and reply to me in the form of JSON.
+    User: The target is the nearest prominent landmark to me. Answer me a dictionary like color:–, feature: –, size: –, type: –.
+    GPT 4o: color: blue, feature: Steel, glass, size: medium size, type: building.
+  • Instruction Fusion.
+    System Prompt: You are an assistant proficient in text processing.
+You need to help me combine these scattered actions and landmarks into a sentence using words with similar meanings and more appropriate words, making them smooth, fluent, and accurate.
+If the landmarks of adjacent actions are similar or even identical, please use pronouns to refer to them.
+    User: Multiple sub-instructions.
+    GPT 4o: Move forward to a high-rise building with a noticeable logo at the top. Then, slightly turn left and go straight to a futuristic tower with a large spherical structure in the middle.
+Appendix G Data Quality Control
 Data Filter.
@@
 Continue heading straight, passing it $\cdots$”, demonstrating the effectiveness of this post-processing technique.
-At the same time, we built a data inspection platform to provide instructions, action sequences, and corresponding images to the examiners.
-If the instructions and trajectories align, they are considered qualified.
-We randomly select 3K samples from the entire dataset according to data distribution in Sec.
-V-A.
-After manually inspecting these samples, we find that the qualification rate reaches 91%.
-IV Dataset Analysis
-TABLE II: Comparison of different VLN datasets.
-Above the middle dividing line lies the ground-based datasets, while below is the aerial VLN datasets.
+Manually Check.
+At the same time, we built a data inspection platform to provide instructions, action sequences, and corresponding images to human examiners.
+If an instruction describes all the actions and landmarks in a trajectory well, it is considered qualified.
+We randomly select 3K samples from the entire dataset according to the data distribution.
+After manually inspecting these samples, we find that they reach a high qualification rate of 91%.
+There is some ambiguity in the description of some landmarks in the remaining data, making it likely that these landmarks are not easily distinguishable from the surrounding environment.
+However, the examiners consider this not entirely unacceptable.
+In summary, most of the generated data feature good quality for the aerial VLN task.
+Appendix H More Dataset Analyses
+Table 4 shows several commonly used statistical information of different VLN datasets, from which we can see that our dataset features a significantly larger number of trajectories and a more extensive vocabulary, as well as greater environmental diversity.
+In contrast, our average trajectory length and instruction length are relatively short.
+This is intentional, as we believe short- and medium-range instructions are actually more in line with the usage habits of human users.
+This might be more beneficial for the aerial VLN field.
+Table 4: Comparison of different VLN datasets.
 $N_{traj}$: the number of total trajectories.
@@
 Intr Len.
-Action Space $N_{act}$ Environment R2R [1] 7189 3.1K 10.0 29 graph-based 5 Matterport3D RxR [25] 13992 7.0K 14.9 129 graph-based 8 Matterport3D REVERIE [37] 7000 1.6K 10.0 18 graph-based 5 Matterport3D CVDN [45] 7415 4.4K 25.0 34 graph-based 7 Matterport3D TouchDown [7] 9326 5.0K 313.9 90 graph-based 35 Google Street View VLN-CE [23] 4475 4.3K 11.1 19 2 DoF 56 Matterport3D LANI [33] 6000 2.3K 17.3 57 2 DoF 116 CHALET ANDH [12] 6269 3.3K 144.7 89 3 DoF 7 xView AerialVLN [29] 8446 4.5K 661.8 83 4 DoF 204 AirSim + UE CityNav [26] 32637 6.6K 545 26 4 DoF - SensatUrban OpenUAV [47] 12149 10.8K 255 104 6 DoF 264 AirSim + UE Ours 100K 15.6K 99.1 59 4 DoF 35 AirSim + UE, GTA5 + Script Hook V, Google Earth Studio, 3D GS + SIBR viewers
-(a) Difficulty level distribution.
-(b) Length-height distribution.
-(c) Action distribution with 1 type of ‘Forward’.
-(d) Action distribution with 3 types of ‘Forward’.
-Figure 4: Statistical analysis of trajectories.
-IV-A Overview
-Using our toolchain, we collect 100k trajectories from 18 scenes, along with corresponding image sequences and language instructions.
-During the data generation process, we set a minimum motion step size of 3 meters to produce more granular trajectories.
-The details of our and previous VLN datasets are listed in Table.
-II, from which we can see that our dataset features a significantly larger number of trajectories and a more extensive vocabulary, as well as greater environmental diversity.
-In contrast, our average trajectory length and instruction length are relatively short.
-This is intentional, as we believe” short- and medium-range instructions are actually more in line with the usage habits of human users.
-This might be more beneficial for the aerial VLN field.
-IV-B Trajectory Analysis
-In addition to a rich variety of scenes, we also strive for diversity in the difficulty level, length, and height of the trajectory data.
-Based on the number of actions in one trajectory, we classify trajectories with fewer than 30 actions as ‘Easy’, those with the number of actions ranging from 30 to 60 as ‘Moderate’, and those with more than 60 actions as ‘Hard’.
+Action Space $N_{act}$ Environment R2R [3] 7189 3.1K 10.0 29 graph-based 5 Matterport3D RxR [4] 13992 7.0K 14.9 129 graph-based 8 Matterport3D REVERIE [2] 7000 1.6K 10.0 18 graph-based 5 Matterport3D CVDN [5] 7415 4.4K 25.0 34 graph-based 7 Matterport3D TouchDown [1] 9326 5.0K 313.9 90 graph-based 35 Google Street View VLN-CE [6] 4475 4.3K 11.1 19 2 DoF 56 Matterport3D LANI [7] 6000 2.3K 17.3 57 2 DoF 116 CHALET ANDH [20] 6269 3.3K 144.7 89 3 DoF 7 xView AerialVLN [17] 8446 4.5K 661.8 83 4 DoF 204 AirSim + UE CityNav [21] 32637 6.6K 545 26 4 DoF - SensatUrban OpenUAV [18] 12149 10.8K 255 104 6 DoF 264 AirSim + UE Ours 100K 15.6K 99.1 59 4 DoF 35 AirSim + UE, GTA V, Google Earth Studio, 3D GS + SIBR viewers
+Figure 7: The distribution of the data volume in different scenes under the Train and Test sets. (a) Train set distribution. (b) Test set distribution.
+Following previous studies [17, 20], we conducted a statistical analysis of the linguistic phenomena using 25 randomly selected instructions and compared the results with other VLN datasets, as detailed in Table 5.
+The analysis shows that the generated instructions exhibit rich linguistic phenomena such as ‘Reference’ and ‘Comparison’.
+Notably, our dataset is not the most complex one, since we believe that instructions in VLN tasks should be more aligned with real-life scenarios, rather than emphasizing length and complexity.
+This cognition is consistent with that of REVERIE [2].
+Our instructions avoid overly lengthy and unrealistic expressions to some extent, making them more practical to command UAVs.
 Fig.
-4(a) shows the corresponding difficulty level distribution.
-Besides, compared with ground-based VLN, the aerial VLN task has more motion dimensions.
-Therefore, we set different trajectory lengths and flight heights to obtain rich data.
+7(a) shows the data distribution of the train set, where 7 UE scenes account for 75.7% of the total 100K data, 4 3D GS scenes account for nearly 20% of the total amount, and Google Earth data accounts for 4.46%.
 Fig.
-4(b) exhibits the distribution of these data, with their lengths ranging from 0 to 300 meters, and the heights ranging from 0 to 150 meters.
-In the aerial VLN tasks of large-scale outdoor scenes, the proportion of moving forward is naturally higher than that of making adjustments in direction and altitude, as shown in Fig.
-4(c).
-However, this highly unbalanced action distribution might cause the VLN model to overfit to the dominant action.
-To alleviate this problem, we divide the ‘Forward’ action into three granularities, i.e., 3m, 6m, and 9m.
-In the ground-truth trajectories, three consecutive ‘Forward’ actions will be combined into one ‘9m Forward’ action.
-At the end of a straight-moving trajectory, if the remaining distance is less than 9m, it will be combined into a ‘6m Forward’ action, or remain as a ‘3m Forward’ action.
+7(b) presents the data distribution of the test set, where the seen data and unseen data account for 60% and 40%, respectively.
+Table 5: Linguistic phenomena analysis using randomly selected 25 instructions.
+$p$ denotes the proportion of instructions showing the phenomenon, and $\mu$ represents the average number of times the phenomenon occurs in each instruction.
+R2R[3] ANDH[20] AerialVLN[17] OpenFly
+Phenomenon $p$ $\mu$ $p$ $\mu$ $p$ $\mu$ $p$ $\mu$ Example in OpenFly
+Reference 100 3.7 92 1.9 100 9.7 100 2.7 …Advance to the large beige and brown building with windows
+Coreference 32 0.5 8 0.1 68 1.8 52 1.4 …Continue moving forward to reach it
+Comparison 4 0.0 32 0.4 20 0.2 60 0.7 …Move ahead to the medium-sized beige building…
+Sequencing 16 0.2 8 0.1 68 3.7 64 0.8 …move forward to next large light brown building …
+Allocentric Relation 20 0.2 32 0.4 56 4.6 76 1.0 …black large building featuring a billboard on its rooftop…
+Egocentric Relation 80 1.2 32 0.4 100 7.1 100 2.4 … then slightly turn left as you move ahead towards…
+Imperative 100 4.0 100 1.1 100 6.9 100 3.6 …. Proceed slightly straight and turn left ….
+Direction 100 2.8 100 1.4 100 4.6 100 3.8 turn right and go ahead to …
+Temporal Condition 28 0.4 20 0.2 76 5.6 72 0.9 …Continue straight until you reach …
+Appendix I Implementation and Training Details
+[…]
+Appendix J Qualitative Experimental Results
 Fig.
-4(d) presents the action distribution after this action merging process.
-(a) Train set distribution.
-(b) Test set distribution.
-Figure 5: The distribution of the data volume in different scenes under the Train and Test sets.
-V OpenFly-Agent
-Figure 6: The architecture of OpenFly-Agent.
-Keyframes at the time of action transitions are selected to extract crucial observations as the history, with corresponding visual tokens compressed to reduce the computational burden.
-V-A Dataset Split
-Similar to previous works, we divide the dataset into three splits, i.e., Train, Test Seen, Test Unseen.
-Detailed data distributions are shown in Fig.
-5.
-For the Train split, 7 scenes under the UE rendering engine account for $75.7\%$ of the total 100K data, since UE provides the largest number of scenes, where different amounts of trajectories are sampled according to the scenario area.
-The 4 scenes created by 3D GS are also the main part of the data, accounting for nearly $20\%$ of the total amount.
-To ensure visual quality, we only collect data from a high-altitude perspective using Google Earth, which accounts for $4.46\%$.
-The detailed information of the Test Seen and Test Unseen splits are as follows:
-  • Test Seen: 1800 trajectories uniformly sampled from 11 previously seen UE and 3D GS scenarios.
-  • Test Unseen: 1200 trajectories uniformly generated from 3 unseen scenarios, i.e., UE-smallcity, 3D GS-sjtu02, and a Los Angeles-like city in GTA V.
-V-B Problem Definition
-In the aerial VLN task, a UAV is randomly positioned within a 3D environment with its initial pose defined as $P=[x,y,z,\phi,\theta,\psi]$.
-At each timestamp $t$, the UAV perceives the surrounding environment through an egocentric image as its observation.
-Guided by natural language instructions, the task involves predicting the next navigation action.
-Notably, the UAV can utilize either the current observations or the frames from all previous timestamps to make its prediction.
-V-C Model Architecture
-As shown in Fig.
-6, we take OpenVLA [22] as the baseline and design an end-to-end model for aerial VLN.
-In contrast, our model takes a sequence of images to indicate the observation instead of one image in the original OpenVLA.
-Moreover, to mitigate visual redundancy between adjacent video frames while maintaining key information, two strategies are proposed, i.e., keyframe selection and visual token merging.
-First, a series of candidate keyframes are selected.
-Then, these keyframes are merged temporally before and after the vision encoder, resulting in a compact sequence of visual tokens.
-Finally, the action decoder discretizes the predicted tokens into uniformly distributed bins, which are subsequently mapped to the 6 action types specific to drones.
-V-C1 Keyframe Selection
-The length of contextual visual tokens is a major challenge for VLMs when processing videos.
-Many open-source VLMs use uniform frame sampling [5, 39, 48] to reduce calculation, but this strategy is not suitable for aerial VLN, since it may miss frames containing key landmarks.
-To address this issue, we adopt a heuristic method to identify keyframes by detecting the change point of the UAV’s movement.
-We notice that sudden changes in the UAV’s trajectory are often caused by the observation of landmarks, which can serve as cues to determine keyframes.
-Specifically, we use the movement of the drone over time to draw turning curves, and the frames near the peaks of the wave are selected as candidate keyframes.
-The resulting data is interpolated and smoothed, forming a wave-like curve that represents the UAV’s movement.
-To further ensure the precision of training data, scene segmentation maps collected in Sec.
-III-A are used on selected frames to detect key landmarks.
-Frames containing landmarks are selected as keyframes, yielding reasonably accurate results.
-Note that each sudden change of actions, e.g., from ‘Forward’ to ‘Turn Left’, will produce a set of keyframes.
-Consequently, we obtain several sets of keyframes for a long trajectory.
-V-C2 Visual Token Merging
-To further reduce redundant information in keyframes, we design visual token merging, where the core concept is to recognize the similarity between image tokens.
-It compares adjacent keyframes to merge similar regions and maintains its simplicity by token compression.
-TABLE III: Comparison results on the test-seen split.
-[…]
-For each set of candidate keyframes obtained in the previous selection process, a visual encoder maps each input image to multiple visual tokens, with each token representing the information of an image patch.
-Considering the potential inter-frame patch redundancy, we take a strategy that similar tokens in subsequent adjacent frames are periodically merged.
-Specifically, we select the first frame in a keyframe set as the reference, since it usually contains the crucial observation indicating the time for action transition.
-Then, we densely calculate the cosine similarities between each pair of visual tokens of the reference image and the subsequent image.
-Next, we merge the tokens with high similarity by averaging them.
-The unmerged tokens in the subsequent frame will be discarded.
-The merging operation is iteratively performed until the entire keyframe set has been traversed.
-Besides, we maintain a memory bank with a capacity of $K$ images, which follows a first-in-first-out (FIFO) policy to retain the latest keyframes.
-After the above process, $M$ visual tokens $E=\{e_{1},e_{2},\cdots,e_{M}\}$ are obtained for each set of keyframes.
-Since aerial VLN requires UAVs to perform long-distance flights based on instructions, we continue to carry out token compress to reduce the computational burden.
-The compressed visual tokens $E_{c}$ are obtained through grid pooling [27].
-Notably, we keep the visual tokens of the current frame uncompressed to capture the latest visual observation, as it contains the most important information for flight action prediction.
-V-C3 Action Prediction
-[…]
-TABLE IV: Comparison results on the test-unseen split.
-[…]
-TABLE V: Ablation study on the test-seen split. ‘KS’ and ‘VTM’ denote keyframe selection and visual token merging, respectively.
-[…]
-VI Experiments
-VI-A Implementation and Training Details
-The proposed OpenFly-Agent is composed of the Dino-SigCLIP (224 $\times$ 224 pixels) as a vision encoder and the pretrained Llama-2 (7B) as a language model.
-The visual tokens extracted by Dino-SigLIP are aligned to the same dimension as text embeddings using a projection layer, and then fed into the Llama-2 model.
-The current frame during flight remains 256 tokens and all historical keyframes are compressed into 1 token.
-The capacity $K$ of the history memory bank is set to 2 in our experiment.
-For action prediction, the last 256 tokens in the vocabulary are used as special tokens for action representation.
-VI-B Evaluation Metrics
-Four standard metrics in VLN tasks are adopted to evaluate different methods, i.e., navigation error (NE), success rate (SR), oracle success rate (OSR), and success weighted by path length (SPL).
-NE measures the average deviation between the UAV’s final stopping point and the ground-truth destination.
-SR calculates the proportion of successful tasks, where a task is considered successful if the UAV stops within 20 m of the target.
-In OSR, if any point on the trajectory is within 20 m of the target, the task can be considered successful.
-SPL calculates the success rate weighted by the ratio of the ground-truth path length to the actually-executed path length.
-VI-C Results
-VI-C1 Quantitative results
-We test the proposed OpenFly-Agent on the test-seen and test-unseen splits and differentiate among easy, moderate, and hard difficulty levels.
-Multiple VLN methods are compared with ours, where ‘Random’ means an agent randomly selects one action to execute until the ‘stop’ action is chosen, Seq2Seq [23] adopts a simple recurrent model for action sequence prediction, CMA [23] uses a bi-directional LSTM to implement cross-modal attention between instructions and images, AerialVLN [29] combines CMA and a look-ahead guidance strategy to enhance robustness.
-Navid [56] leverages the strong ability of VLMs and develops a video-based architecture, achieving impressive results in indoor VLN.
-Table.
-III shows quantitative results on the test-seen split.
-Seq2Seq, CMA, and AerialVLN achieve limited success rates.
-Counterintuitively, they perform worse on easy samples.
-We find that their OSR is not low while they are difficult to stop at a short-range distance.
-In contrast, Navid obtains better results, demonstrating the great potential of VLMs in aerial VLN.
-Our OpenFly-Agent achieves the best performance for its strong VLM backbone and the proposed strategies, while there is still much room for improvement.
-Table.
-IV presents quantitative results on the test-unseen split, illustrating the generalization capabilities of these VLN methods.
-Similarly, our method achieves better performance, exhibiting a certain degree of robustness.
-However, all methods are significantly degraded, which indicates models with greater generalization are urgently needed to be developed.
-Figure 7: Illustration of an aerial VLN trajectory generated by OpenFly-Agent, which successfully predicts actions following the instruction when encountering landmarks.
+8 presents a qualitative result in a UE scene, where our OpenFly-Agent successfully navigates to the destination according to the instruction.
+It presents a powerful capability in perceiving environments and aligning observations with complex instructions.
+Fig.
+9 presents another successful aerial VLN example in a 3D GS scene.
+The image style, flight heights, and viewpoints are significantly different from UE’s scenarios.
+In this case, our OpenFly-Agent exhibits robustness to handle data with great diversity.
+In addition, Fig.
+10 shows two failure cases, where our model sometimes fails to identify the landmark or output actions with proper amplitudes.
+Figure 8: Illustration of an aerial VLN trajectory generated by OpenFly-Agent in a UE scene, which successfully predicts actions following the instruction when encountering landmarks.
+Table 6: Results of real-world experiments.
+model SR $\uparrow$ OSR $\uparrow$
+Navid [16] 13.0% 17.39%
+NaVILA [43] 13.0% 21.74%
+OpenFly-Agent (Ours) 26.09% 34.78%
+Table 7: Performance of OpenFly-Agent trained on different datasets.
+dataset SR $\uparrow$ OSR $\uparrow$
+AerialVLN  [17] 4.3% 13.04%
+OpenFly (Ours) 26.09% 34.78%
+Figure 9: Illustration of aerial VLN trajectories generated by OpenFly-Agent in a 3D GS scene.
 Figure 10: Illustration of failure cases. Sometimes our model may misclassify key landmarks or output wrong actions.
-VI-C2 Qualitative results
-Fig.
-7 presents a qualitative result, where our OpenFly-Agent successfully navigates to the destination according to the instruction.
-It presents a powerful capability in perceiving environments and aligning observations with complex instructions.
-In addition, Fig.
-8 shows two failure cases, where our model fails to identify the landmark or output actions with proper amplitudes.
-VI-D Ablation Study
-Ablation studies are conducted to evaluate the contribution of the keyframe selection and visual token merging in OpenFly-Agent.
-Table.
-V shows the results, where OpenVLA [22] is our baseline.
-Using only the current frame as an observation makes OpenVLA perform poorly in the aerial VLN task.
-From ‘History + VTM’ we can see that historical frames significantly improve the success rate.
-The keyframe selection strategy further increases the SR from 12.7% to 18.5%, demonstrating the effectiveness of key observations.
-Besides, the comparison between ‘KS’ and ‘KS + VTM’ indicates the great effect of our visual token merging strategy.
-VII Limitations and Future Work
-The proposed OpenFly platform incorporates various rendering engines/techniques to provide high-quality scenes.
-Specifically, this is the first attempt to use 3D GS reconstructed scenes to support real-to-sim training and testing, while in the reconstruction of large-scale areas, a few visual artifacts are inevitably present.
-Future work will focus on exploring more effective reconstruction methods to enhance realism in large-scale scenes.
-Besides, the proposed OpenFly-Agent is built upon the large VLN model architecture, which is not practical for real-time deployment on UAVs.
-To address this, future research should focus on developing more efficient architectures and effective quantization techniques.
-VIII Conclusion
-In this work, we present OpenFly, a platform designed for large-scale data collection in aerial Vision-and-Language Navigation (VLN).
-OpenFly integrates multiple rendering engines and advanced real-to-sim techniques for data generation, enabling efficient collection of diverse, high-quality aerial VLN data.
-The resulting large-scale dataset comprises 100k trajectories across 18 distinct scenes, spanning a wide range of altitudes and difficulty levels, which is significantly superior than existing ones.
-Furthermore, we propose OpenFly-Agent, a keyframe-aware aerial navigation model capable of directly predicting flight actions based on observations and language instructions.
-Extensive experiments validate the effectiveness of the proposed method, and establishing a comprehensive benchmark for future advancements in aerial navigation.
+Appendix K Real-world Experiments
+The real-world experiments are conducted in 23 real outdoor scenes, where each scene corresponds to an unseen VLN task created by human operators, and the trajectory lengths range from 50m to 500m.
+We use a Q250 airframe as a real agent, carrying an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as the onboard computer.
+All methods run on an external PC communicating with the onboard computer to transfer images and action instructions.
+Two most recent models, Navid [16] and NaVila [43], are evaluated for comparison.
+The results are shown in Table 6, where our model achieves the best performance with 26.09% SR and 34.78% OSR, significantly outperforming the comparison methods.
+This experiment again indicates the superiority of our OpenFly-Agent.
+Besides, we also trained our model on both our own dataset and the AerialVLN dataset separately.
+The results are shown in Table 7, where the first row is trained on the AerialVLN dataset (without real-to-sim data), and the second row is trained on our OpenFly dataset.
+It strongly demonstrates the capability of our data generation method in bridging the sim-to-real gap.
+A qualitative result is presented in Fig.
+11, and a dynamic demo can be found in our supplementary video.
+Figure 11: Snapshots of the real-world experiment.